# Privacy in Speech Technology

### v2

@@ IV-A3 Disentanglement
IV-B Secure processing
If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key. Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [yao1986generate, portelo2015logsum, portelo2014privacy, Teixeira2018, cramer2015mpc] (see fig. 15). The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, and garbled circuits. The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task. Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved. Another category of approaches is distance-preserving hashing techniques, where data is obfuscated through randomization such that the desired communication is preserved while other private information is kept secret [portelo2015privacy, jimenez2015secure].
Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [thaine19_interspeech, Zhang2019]. It can also be applied in privacy-preserving speaker recognition, authentication or verification [teixeira22_interspeech, pathak11_interspeech, pathak2012ppsa, portelo2014bprivacy], speech search [portelo2015bprivacy], or for secure sound classification [shashanka2006secure]. Finally, we can compare the similarity of audio signals without revealing them [rane2013privacy, perez2020, zarazaga2022thesis].
A central problem with such cryptographic protocols is that the operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [smaragdis2007framework, pathak2012ppsv, pathak2011privacy, pathak2013privacy]. In contrast, typical speech-processing algorithms also require other operations, such as non-linearities, conditionals, and recursions [chialva2018conditionals], making the algorithms impossible to implement. While many operations have standard workarounds [teixeira2019privacy], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g. [thaine19_interspeech]. In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large. For a broader overview of cryptographic methods in speech, see [nautsch2019preserving].
@@ IV-B Secure processing
Figure 17: Federated learning as an example of distributed learning, where devices train models locally and transmit only model updates to the cloud. Red dashed arrows marked with exclamation marks indicate the reduced flow of information.
IV-C Privacy-preserving architectures
@@
Perhaps the most obvious solution to protecting privacy, for example in the Cloud Leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig. 16). Such limitation or reduction in information flow can be implemented by removing, replacing, or distorting private content as in fig. 16(a) (cf. section IV-A). At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig. 16(b) [shi2016edge]. In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked. Observe, for protecting privacy in scenarios 1 to 4 in fig. 1, local processing is the best protection as nothing is transmitted to a cloud service.
A majority of advanced speech processing today, however, uses machine learning, which has to be trained using large databases of speech. The best quality data correspond closely to scenarios where the services are used. Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case. This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
@@
Distributed learning is an approach to training models without the need for centralized data collection [liu2021machine]. Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig. 17). The model updates are, however, calculated from the raw signal and thus describe the original signal in some convoluted way. Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf. section IV-A).
@@
Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates. It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [feng22b_interspeech, granqvist20_interspeech, zewoudie2021federated, nelus2021, shoemate2022sottovoce, ro21_interspeech, nguyen2023federated, feng2021attribute].
@@
Overall, distributed learning is a promising approach, even if it has several challenges. First, constructing, training, and testing systems architectures is much more complicated than regular machine learning. Second, model updates are model-specific and cannot easily be reused if the model structure is updated. The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches. Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [Orekondy2018gradientleaks, liu2021machine]. To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig. 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
IV-D Acoustic interventions
@@
When a user interacts with a speech interface, then the spoken answer of the speech interface can contain private information. This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs. 19 and III-D5). In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig. 18). However, tracking itself also poses a threat to privacy, as it is yet another form of private information extracted from the users. This places great trust in the local device’s ability to track and identify the authorization levels of the people present in the acoustic space. Such systems have, however, not yet been widely published.
@@
Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [gardner1998_3daudio, pulkki2018parametric]. This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g. [donley2016, nielsen2018, WALLACE2022101285]) or as an audio spotlight through non-linear interference effects between ultrasound sources [yoneyama1983audio]. By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
@@
Note that sound zones with destructive interference do receive a partial observation of the private message. Still, it is (hopefully) low-energy and distorted to the extent that it is unintelligible. The central challenges of this approach are to make the constructive sound zone large enough that it allows for small head movements and to make the destructive interference uniform everywhere else, such that there are no isolated points with constructive interference outside the desired sound zone. A benefit of this approach is, however, that it requires tracking of only the target listener, which is, while difficult, still much simpler than tracking all the people in the room.
IV-E Improving Performance
@@
In many use cases, like False Activation, the actual culprit is the inadequate performance of the service. For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy. Since the inadequate wake word detector is thus causing a privacy breach, the best solution is to improve the wake word detector. The treatment thus addresses the cause rather than the symptoms thereof. Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [Kalloniatis2009, van2011privacy].
@@
This approach has, however, two principal challenges. First, improving performance often requires an increase in computational power and other resource consumption. This is financially costly and has an environmental penalty [patterson2021carbon]. In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used. Second, speech interfaces will always have occasional errors, even with improved performance. That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [hong2004privacy, backstrom2021edpbcomments]. Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar. The intrusiveness of such additional protections should then reflect the severity of the potential breach such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
@@
User Alice Local device Noise Remote service Refer to caption Refer to caption $+$ Refer to caption !
@@
TABLE III: Dimensions of performance in privacy-preserving speech processing.
@@
V Evaluating the Performance of Privacy-Preserving Methods
To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig. 2). With utility, we refer here to the usefulness of the service in its intended task. While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good. Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table III). Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections V-A, V-B, V-C and V-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VI). We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-E).
The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [tomashenko2022voiceprivacy].
V-A Objective measures of privacy
Evaluation of objective privacy can be approached through two alternative approaches, through empirical methods or using theoretical concepts. The two categories emanate from different traditions and have their own benefits and areas of application, though the methodologies have some overlaps. More complete integration of the two categories would be highly beneficial, potentially providing the benefits of both categories, but at the time of writing, methods for such unification are not available. More comprehensive treatments of objective privacy metrics are provided in [nautsch2019preserving, noe2022towards].
V-A1 Empirical Privacy
Evaluation of privacy in, for example, the VoicePrivacy challenge [tomashenko2022voiceprivacy], is based on direct measurement of identifiability using signal processing and machine learning methods. The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig. 3).
Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [natgunanathan2016protection]:
@@ V-A1 Empirical Privacy
    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [maouche20_interspeech, ISObiometric]. An increase in EER means that the attacker has made more errors, and privacy has improved.
@@
    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [maouche20_interspeech, brummer2006application].
  5. 5.
    The expected privacy disclosure from the Privacy ZEBRA framework [nautsch20_interspeech] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
  6. 6.
    The worst-case disclosure also from the Privacy ZEBRA framework [nautsch20_interspeech] is the largest disclosure observed in a dataset.
Analysis of anonymized speech signals is straightforward since the above measures can be applied with conventional speech analysis methods to identify the anonymized attributes. Existing methodologies for evaluating speech analysis algorithms can thus be reused such that the evaluation is based on a body of prior research. On the other hand, empirical evaluation of any other form of signal, such as embeddings of machine learning representations or the compressed data of a telecommunications channel, requires analysis methods developed specifically for that data format. While the benefit is that such analysis directly evaluates the signal with any potentially confounding or obfuscation due to additional processing, the overhead of purpose-built analysis can be great.
A central downside of empirical privacy evaluation is that it is based on the current state-of-the-art, which can be surpassed by future technology. In other words, weaknesses in anonymization that are not (yet) known to the evaluator can compromise privacy. Moreover, it is widely known that the performance of machine learning models can be improved by making the model larger or using more data to train such models. A speech analysis model can thus be (likely) improved, not only by future developments but with this currently known methodology that only requires additional effort. Empirical evaluation can thus never give proofs, but just characterizations of the level of privacy.
As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [tomashenko2022voiceprivacy]. Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [KINNUNEN201012]. The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [vali2024porcupine].
V-A2 Theoretical Privacy
Theoretical evaluation of privacy aims to give assurances or proofs of the level of privacy based on notions of information theory. The idea is to quantify the new information about a speaker gained through an observed speech signal in terms of entropy and units of nits or bits.
In other words, suppose an attacker has prior assumptions about a speaker, quantified as a probability distribution $p(X)$ of the possible states $X$ of the speaker. The information content of that model is the negative logarithm $I(X)=-\log p(X)$ and its entropy the expectation thereof $H(X)=E\left[I(X)\right]$. The attacker then gains access to or makes an observation $Y$. This new information modifies the attacker’s beliefs about the speaker, expressed as the conditional model of the speaker $X$ given the information $Y$, or $X|Y$. Correspondingly, the posterior, conditional entropy $H(X|Y)$ quantifies the amount of information the attacker has about the speaker after (posterior) to the observation.
$\epsilon=H\left(X\right)-H\left(X|Y\right).$ (1)
The leak is thus equal to the mutual information of $X$ and $Y$.
A related concept is the log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [nist2021SRE]
@@ V-A2 Theoretical Privacy
Another important and related formulation is differential privacy, which quantifies how the amount of information about a speaker’s affiliation to a particular group changes due to an observation [dwork2006differential, dwork2008differential, nautsch2019preserving]. In particular, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$. The leakage of information is then new information gained through a slightly different dataset $D^{\prime}$ as
Leaks can also be formulated as a function of the groups’ sizes where the speaker is known to belong before and after an observation. Suppose we know that a speaker $x$ belongs to group $x\in A=\{a_{1},\dotsc,\,a_{M}\}$ with $M$ members. The speaker can be any one of the group members with equal likelihood $p\left(x=a_{k}\right)=\frac{1}{M}$. The identity within this group, which is unknown to the attacker, then has an information content of $\log M$ nits. If an attacker somehow manages to identify that the speaker belongs to a subset $B\subset A$ of size $N$, then the remaining information content unknown to the attacker is $\log N$ nits. The reduction in uncertainty is thus the leak magnitude
$\epsilon=\log M-\log N=\log\frac{M}{N}.$ (4)
A leak is thus a function of relative group size before and after an observation.
It is arguably more informative for users to state the absolute group size within which they remain anonymous. If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [sweeney2002k].
A central issue with theoretical metrics is that they are based on mathematical definitions, whereas speech is a natural signal that does not submit to exact formulas [nautsch20_interspeech]. To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters. Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [shoemate2022sottovoce], emotion recognition [feng22b_interspeech], speaker anonymization [shamsabadi2023diff], and federated learning in acoustic sensor network [el2022differential]. As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
V-B Subjective evaluation of privacy

### v3

@@ V-A3 Disentanglement
V-B Secure processing
If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key. Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [131, 132, 133, 108, 134] (see fig. 15). This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, secret sharing, and garbled circuits. The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task. Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved. Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [21, 135].
Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [136, 137]. It can also be applied in privacy-preserving speaker recognition, authentication or verification [31, 39, 41, 43], speech search [138], or for secure sound classification [139]. Finally, we can compare the similarity of audio signals without revealing them [140, 141, 18].
A central problem with such cryptographic protocols is that the floating- or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [142, 143, 144, 145]. The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [146]. In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [147], making the algorithms impossible to implement. While many operations have standard workarounds [148], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g. [136]. In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large. For a broader overview of cryptographic methods in speech, see [13].
@@ V-B Secure processing
Figure 17: Federated learning as an example of distributed learning, where devices train models locally and transmit only model updates to the cloud. The red dashed arrows marked with exclamation marks indicate the reduced flow of information.
V-C Privacy-preserving architectures
@@
Perhaps the most obvious solution to protecting privacy, for example in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig. 16). Such limitation or reduction in information flow can be implemented by removing, replacing, or distorting private content as in fig. 16(a) (cf. section V-A). At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig. 16(b) [149]. In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked. Observe, for protecting privacy in scenarios 1 to 4 in fig. 1, local processing is the best protection as nothing is transmitted to a cloud service.
@@
Distributed learning is an approach to training models without the need for centralized data collection [150]. Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig. 17). The model updates are, however, calculated from the raw signal and thus describe the original signal in some convoluted way. Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf. section V-A).
@@
Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates. It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [91, 37, 29, 80, 59, 151, 60, 152].
@@
Overall, distributed learning is a promising approach, even if it has several challenges. First, constructing, training, and testing systems architectures is much more complicated than regular machine learning. Second, model updates are model-specific and cannot easily be reused if the model structure is updated. The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches. Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [153, 150]. To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig. 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
V-D Acoustic interventions
@@
When a user interacts with a speech interface, then the spoken answer of the speech interface can contain private information. This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs. 19 and IV-D5), as in attack vectors False activation (section IV-D2), WASN authentication (section IV-D4), Speech interface leak (section IV-D5), Discussion leak (section IV-D6), and Shared device (section IV-D7). In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig. 18). However, tracking itself also poses a threat to privacy, as it is yet another form of private information extracted from the users. This places great trust in the local device’s ability to track and identify the authorization levels of the people present in the acoustic space. Such systems have, however, not yet been widely published.
@@
Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [154, 155]. This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g. [156, 157, 158]) or as an audio spotlight through non-linear interference effects between ultrasound sources [159]. By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
@@
Note that sound zones with destructive interference do receive a partial observation of the target information. Still, it is (hopefully) low-energy and distorted to the extent that it is unintelligible. The central challenges of this approach are to make the constructive sound zone large enough that it allows for small head movements and to make the destructive interference uniform everywhere else, such that there are no isolated points with constructive interference outside the desired sound zone. A benefit of this approach is, however, that it requires tracking of only the target listener, which is, while difficult, still much simpler than tracking all the people in the room.
V-E Improving utility
@@
In many use cases, like False activation (section IV-D2), Speech interface leak (section IV-D4), and Discussion leak (section IV-D6), the actual culprit is the inadequate performance of the service. For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy. Since the inadequate wake word detector is thus causing a privacy breach, the best solution is to improve the wake word detector. The treatment thus addresses the cause rather than the symptoms thereof. Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [160, 161].
@@
This approach has, however, two principal challenges. First, improving utility often requires an increase in computational power and other resource consumption. This is financially costly and has an environmental penalty [162]. In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used. Second, speech interfaces will always have occasional errors, even with improved utility. That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [163, 164]. Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar. The intrusiveness of such additional protections should then reflect the severity of the potential breach such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
@@
User Alice Local device Noise Remote service Refer to caption Refer to caption ++ Refer to caption !
@@
V-F Privacy-by-design and provable protections
Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [165]. This increases the chances that the system does not violate privacy unnecessarily, since the privacy-preserving performance is considered at every step of the design process.
The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [166, 167, 13]. Privacy is generally enforced by reducing the accuracy of private information sufficiently, for example, through the addition of noise or by quantization (see fig. 20 and cf. section V-A). To determine how much accuracy must be reduced, differential privacy quantifies how the amount of information available about a speaker’s affiliation to a particular group changes due to an observation. Specifically, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$. The limit $\epsilon$ for leakage of information (in nits) is then the amount of new information gained through a slightly different dataset $D^{\prime}$ as
Leaks can also be formulated as a function of the groups’ sizes, where the speaker is known to belong before and after an observation. Suppose we know that a speaker $x$ belongs to group $x\in A=\{a_{1},\dotsc,\,a_{M}\}$ with $M$ members. Without prior knowledge, the speaker can be any one of the group members with equal likelihood $p\left(x=a_{k}\right)=\frac{1}{M}$. The identity within this group, which is unknown to the attacker, then has an information content of $\log M$ nits. If an attacker somehow manages to identify that the speaker belongs to a subset $B\subset A$ of size $N$, then the remaining information content unknown to the attacker is $\log N$ nits. The reduction in uncertainty is thus the leak magnitude
$\epsilon=\log M-\log N=\log\frac{M}{N}.$ (2)
A leak is thus a function of the relative group size before and after an observation.
It is arguably more informative for users to state the absolute group size within which they remain anonymous. If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [168].
Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy. Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level. Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity. New samples are compared to such biometric references to find matching attributes. The three key concepts in such evaluation are [97, 13]
  • Renewability – multiple, unlinkable biometric references can be created from the same biometric samples and used to identify the person without revealing information about the biometric samples.
Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed. At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them. Such concepts are central to the legal definitions of privacy and first proposals for their mathematical characterization have emerged [169, 170].
These theoretical design tools and concepts can be used in the design of, say, information bottlenecks (section V-A), to quantify the maximal leak of private information. However, care must be taken to ensure that prior assumptions, such as the equal likelihood of all labels, are met (or enforced) in practical scenarios (cf. [123]). Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [59], emotion recognition [91], speaker anonymization [53], and federated learning in acoustic sensor network [171]. As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
TABLE IV: Dimensions of performance in privacy-preserving speech processing.
@@ V-F Privacy-by-design and provable protections
VI Evaluating the performance of privacy-preserving methods
To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig. 2). With utility, we refer here to the usefulness of the service in its intended task. While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good. Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table IV). Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII). We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-F).
The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [98]. It is based on direct measurement of identifiability using signal processing and machine learning methods. The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig. 3).
VI-A Objective measures of privacy
An information-theoretic approach to the objective measurement is to quantify the new information about a speaker gained through an observed speech signal in terms of entropy and units of nits or bits. In other words, suppose an attacker has prior assumptions about a speaker, quantified as a probability distribution $p(X)$ of the possible states $X$ of the speaker. The information content of that model is the negative logarithm $I(X)=-\log p(X)$ and its entropy is the expectation thereof $H(X)=E\left[I(X)\right]$. The attacker then gains access to or makes an observation $Y$. This new information modifies the attacker’s beliefs about the speaker, expressed as the conditional model of the speaker $X$ given the information $Y$, or $X|Y$. Correspondingly, the posterior, conditional entropy $H(X|Y)$ quantifies the amount of information the attacker has about the speaker after (posterior to) the observation.
$\epsilon=H\left(X\right)-H\left(X|Y\right).$ (3)
The leak is thus equal to the mutual information of $X$ and $Y$.
A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [87]. To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters. Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [172]:
@@ VI-A Objective measures of privacy
    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [56, 88]. An increase in EER means that the attacker has made more errors, and privacy has improved.
@@
    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [173]
@@
  5. 5.
    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [56, 174].
  6. 6.
    The expected privacy disclosure from the Privacy ZEBRA framework [87] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
  7. 7.
    The worst-case disclosure also from the Privacy ZEBRA framework [87] is the largest disclosure observed in a dataset.
Private information can be potentially extracted (and privacy-preserving performance evaluated) from either 1. the waveform of a speech signal or 2. from any intermediate representation of speech or its attributes within a speech processing system, such as an information bottleneck. In the analysis of waveforms, we can use generic speech analysis tools to determine the magnitude of disclosure. In contrast, intermediate representations depend on the configuration and design of the speech processing system. Analysis and evaluation tools thus have to be purpose-built and/or optimized for every different speech processing system, which is usually a time and labour-intensive effort. The benefit is, however, that when analyzing the intermediate representations, such as an information bottleneck, we analyze a compact representation of the potential leak, without the potentially confounding effects introduced by later processing steps, such as the decoder of an autoencoder. Analyzing waveforms can thus be easier, while analysis of intermediate representations is more effort but potentially more informative.
A central downside of empirical privacy evaluation, such as the analysis of the accuracy of a speaker recognition model, is that it is based on a current state-of-the-art model, which can be surpassed by future technology. In other words, weaknesses in anonymization that are not (yet) known to the evaluator can compromise privacy. Moreover, it is widely known that the performance of machine learning models can be improved by making the model larger or using more data to train such models. A speech analysis model can thus be (likely) improved, not only by future developments but also with this currently known methodology that only requires additional effort. Empirical evaluation can thus never give proofs, but just characterizations of the level of privacy. In contrast, provable privacy can be provided by theoretical frameworks such as differential privacy (cf. section V-F).
As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [98]. Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [35]. The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [123]. More comprehensive treatments of objective privacy metrics are provided in [13, 175].
VI-B Subjective evaluation of privacy